OpenStudy (anonymous):

A hard candy has a radius of 9 mm. The candy has a spherical soft center with a radius of 6 mm, with a hard shell surrounding it. What is the volume of the hard shell? Use 3.14 to approximate pi and express your final answer in hundredths.

OpenStudy (anonymous):

Find the volumes of the two spheres.

OpenStudy (anonymous):

V = 4/3 * pi * r^3 V = 4/3 * 3.14 * 9^3 V = 4/3 * pi * r^3 V = 4/3 * 3.14 * 6^3

OpenStudy (anonymous):

Simplify and subtract them.

OpenStudy (anonymous):

2147.76?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

That's our answer.